Determine the total hardness of a water sample. Given



Volume of EDTA consumed for 20ml of standard hard water (V1) = 17.3 ml




Volume of EDTA consumed for 20ml of the sample hard water (V2) = 15.8 ml

1
Expert's answer
2022-01-10T13:56:31-0500

20-17.3=2.7

20-15.8= 4.2

= 2.7+4.2= 6.9